# Yong Ji1, Junye Li1 and Rui Yang∗2,3

### v3

@@
Abstract. Saturated sets and its reduced case, the set of generic points, constitute two significant types of fractal-like sets in multifractal analysis of dynamical systems. In the context of infinite entropy systems, we aim to provide some qualitative aspects of saturated sets and the set of generic points from both topological and measure-theoretic perspectives.
For systems with the specification property, we establish certain variational principles for saturated sets in terms of Bowen and packing metric mean dimensions, and show the upper capacity metric mean dimension of saturated sets has full metric mean dimension. As applications, we further present some qualitative aspects of the metric mean dimensions of level sets and the set of mean Li-Yorke pairs in infinite entropy systems.
Keywords and phrases: Metric mean dimension; Saturated set; Generic point; Variational principle; Multifractal analysis
@@ 1 Introduction
Let $(X,d,f)$ be a topological dynamical system (t.d.s. for short), where $(X,d)$ is a compact metric space with a metric $d$, and $f:X\to X$ is a homeomorphism. Let $M(X)$, $M_{f}(X)$, $M_{f}^{e}(X)$ denote the sets of Borel probability measures endowed with weak-topology, $f$-invariant Borel probability measures, and ergodic $f$-invariant Borel probability measures on $X$, respectively.
Quantitative analysis using some entropy-like quantities is of great significance in understanding the dynamics of abstract dynamical systems. Measure-theoretic entropy and topological entropy are two crucial invariants that helps us understand the topological complexity of dynamical systems, which are related by the classical variational principle:
@@
where $h_{top}(f)$ is the topological entropy of $X$, $h_{\mu}(f)$ is the measure-theoretic entropy of $\mu$, and the supremum ranges over all $f$-invariant measures of $X$. In 1973, resembling the definition of Hausdorff dimension Bowen [Bow73] introduced Bowen topological entropy for any subset of the phase space $X$. The pioneering work has directly contributed to the development of dimension theory in dynamical systems, especially serving as a powerful tool for studying the “size” of fractal-like sets in multifractal analysis. The multifractal analysis, with its origin dating back to Besicovitch, takes into account the decomposition of the whole phase space into smaller subsets consisting of points with the similar dynamical behaviors(or known as multifractal spectrum), and then, from a geometrical or topological perspective, describes the “size” of those subsets by using fractal-like dimensions and dimensional-like entropies. Such a procedure allows us to get more information about the dynamics and even recovers the partial dynamical behaviors of the system, a phenomenon known as multifractal rigidity.
We aim to study the multifractal analysis of saturated sets in infinite entropy systems. Let us mention some relevant results about the dimensional entropies of saturated sets to clarify the motivation. Given $x\in X$, by
$\mathcal{E}_{n}(x):=\frac{1}{n}\sum_{i=0}^{n-1}\delta_{f^{i}x}\in M(X),$
where $\delta_{x}$ is the Dirac mass at $x$, we denote the $n$-th empirical measure of $x$. Then, in the sense of weak-topology, the limit-point set $V_{f}(x)$ of the sequence $\{\mathcal{E}_{n}(x)\}_{n\geq 1}$ is a nonempty compact connected subset of $M_{f}(X)$ (cf. [DGS76, Proposition 3.8]). Specially, when $V_{f}(x)=\{\mu\}$ for some invariant measure $\mu$, then $x$ is called a generic point of $\mu$. Denote by $G_{\mu}$ the set of all generic points of $\mu$. The Birkhoff’s ergodic theorem implies that $\mu(G_{\mu})=1$ if and only if $\mu$ is ergodic. In 1973, Bowen [Bow73] proved that for any ergodic measure $\mu$, the Bowen topological entropy of $G_{\mu}$ equals $h_{\mu}(f)$, i.e.,
@@
For the case of non-ergodic measure $\mu$, one has $\mu(G_{\mu})=0$, and there even exist the examples showing that $G_{\mu}=\emptyset$ but $h_{\mu}(f)>0$. Therefore, Bowen’s result fails for non-ergodic measures. It is natural to ask whether there are some proper conditions such that the Bowen’s equality still holds for non-ergodic measures.
A positive answer to the above question is due to Pfister and Sullivan [PS07], who provided it within a more general framework. A subset $K$ of $X$ is called a saturated set if $x\in K$ and $V_{f}(x)=V_{f}(y)$, then $y\in K$. More generally, given a nonempty closed connected subset $K\subset M_{f}(X)$, we have a saturated set by defining $G_{K}=\{x\in X:V_{f}(x)=K\}$ associated with $K$. For any topological dynamical system (t.d.s) possessing the specification property, Sigmund [Sig72, Sig74] firstly proved the existence of the saturated sets, and then was later extended to non-uniformly hyperbolic and non-uniformly expanding systems by several authors [LST13, TV17]. In 2007, Pfister and Sullivan [PS07] introduced the $g$-almost product property that is strictly weaker than specification property, which solely requires a point partially $\epsilon$-shadowing the given finite pieces of orbit-segments and the uniform separation property (e.g. expansive and more generally asymptotically $h$-expansive systems), and established the variational principle of Bowen topological entropy of nonempty compact connected set $K\subset M_{f}(X)$:
$h_{top}^{B}(T,G_{K})=\inf\{h_{\mu}(f):\mu\in K\}.$
Furthermore, under only the assumption of $g$-almost product property, they [PS07] showed that for all $\mu\in M_{f}(X)$,
$h_{top}^{B}(f,G_{\mu})=h_{\mu}(f),$
which extends Bowen’s remarkable result to non-ergodic measures. Consequently, the work of [Bow73, PS07] leads to an intensive study toward the estimation of different types of dimensional-like entropies of saturated sets. Under the specification property and positive expansive property, Zhou, Chen and Cheng [ZCC12] also obtained the variational principle of packing topological entropy of nonempty compact connected set $K\subset M_{f}(X)$,
$h_{top}^{P}(f,G_{K})=\sup\{h_{\mu}(f):\mu\in K\}.$
By weakening the specification property and considering systems with the $g$-almost product property, Hou, Tian and Zhang [HTZ23] not only demonstrated that the variational principle $h_{top}^{P}(f,G_{K}^{C})=\sup\{h_{\mu}(f):\mu\in K\}$ holds for (mixed) saturated sets with more complicated dynamics, but also showed that the upper capacity topological entropy of the saturated sets has full topological entropy, i.e.,
@@
where $G_{K}^{C}:=G_{K}\cap\{x\in X:C_{f}(X)\subset\omega_{f}(x)\},$ $\omega_{f}(x):=\bigcap_{n=0}^{\infty}\overline{\bigcup_{k=n}^{\infty}\{f^{k}x\}}$ is the $\omega$-limit set of $x$, and $C_{f}(X)$ is the measure center given by $C_{f}(X):=\overline{\cup_{\mu\in M_{f}(X)}{\rm supp}{(\mu)}}$, ${\rm supp}(\mu)$ is the supported set of $\mu$. Additionally, Huang, Tian and Wang [HTW19] showed that Pfister and Sullivan’s variational principle is still valid for transitive-saturated set $G_{K}^{T}:=G_{K}\cap{\rm Tran}$, where ${\rm Tran}$ means the set of transitive points. The generalizations of the aforementioned results can be found in [JCZ22, LLS17] for the scenarios of nonuniformly hyperbolic systems, and in [ZC18, DZZ23, Z18, RTZ23, ZRZ23] for the cases of the actions of amenable groups
Based on the possible values of topological entropy, dynamical systems can be can be classified into three types: zero entropy systems, finite positive entropy systems and infinite entropy systems. Infinite entropy systems typically imply that these systems may exhibit rather complicated topological complexity. As a result, the concept of topological entropy fails to offer more information about the dynamics of the system. In 2000, Lindenstrauss and Weiss [LW00] introduced a refined notion for topological entropy, called metric mean dimension, which is a dynamical analogue of box dimension to measure how fast the $\epsilon$-topological entropy diverges to $\infty$ as $\epsilon$ goes to 00. It has been shown that every system with finite entropy has a metric mean dimension of zero. See some applications of metric mean dimension in the analog compression of information theory [GS20], and in estimating the upper bound of mean dimension of some complex systems [Tsu18a, Tsu18b]. The notion of measure-theoretic counterpart of metric mean dimension, called upper and lower mean Rényi information dimensions, is due to Gutman and Śpiewak[GS21]. More precisely, the lower and upper mean Rényi information dimensions of $\mu\in M_{f}(X)$ are respectively defined by
$\underline{\rm MRID}_{\mu}(f,X,d):$ $=\liminf_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{\mathrm{diam}(\xi)<\epsilon}h_{\mu}(f,\xi),$
$\overline{\rm MRID}_{\mu}(f,X,d):$ $=\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{\mathrm{diam}(\xi)<\epsilon}h_{\mu}(f,\xi),$
where the infimum ranges over the finite Borel measurable partition with diameter at most $\epsilon$. This measures how fast the Kolmogorov-Sinai $\epsilon$-entropy $\inf_{\mathrm{diam}(\xi)<\epsilon}h_{\mu}(f,\xi)$ diverges to $\infty$ as $\epsilon\to 0$. Several variational principles regarding the metric mean dimension of $\mathbb{Z}$-actions can be found in the works [LW00, GS21, S22, YCZ23a, YCZ25]. Then, in the context of the infinite entropy systems, the aforementioned work suggests the following questions for saturated sets:
@@
For (1), the authors in [YCZ25] showed that the packing metric mean dimension of the set of generic points is equal to the mean Rényi information dimension. However, this equality holds for Bowen metric mean dimension under the assumption of an analogous Brin-Katok formula within the context of infinite entropy systems (cf.[YCZ25, §5, Question,(1)]); for (2), since the uniform separation property implies the system has finite topological entropy, extending the work of [PS07] to the framework of infinite entropy systems encounters some nontrivial challenges. Under the assumption of $g$-almost product property, for any non-empty compact connected set $K$ contained in the convex closure of finite invariant measures, Yuan [Y23] investigated the Bowen metric mean dimension of $G_{K}$.
Prior to the comparison of the different types of specification-like property, we concentrate solely on the systems possessing the specification property and provide a comprehensive answer to question (2). Actually, since the asymptotic behavior of orbits ${\rm orb}_{f}(x)=\{f^{n}(x):n\geq 0\}$ is a fundamental topic in ergodic theory, the subsequent statements of main results involve more general (mixed) saturated sets. For simplicity, we write
@@
where $U\subset X$ is a nonempty open subset, $K\subset M_{f}(X)$ is a nonempty closed connected (or convex) subset. If $G_{K}^{\omega,U}\not=\emptyset$, then there exists a transitive point in the open set $U$ such that $V_{f}(x)=K$.
Saturated sets exhibit different dynamical behaviors in terms of different types of metric mean dimensions. There are indeed certain variational principles for the Bowen and packing metric mean dimensions of saturated sets, as presented in Theorems 1.1, 1.2 and 1.3. Although the $\mu$-measure of $G_{\mu}$ is zero for every $\mu\in M_{f}(X)$, its topological complexity can be sufficiently large, that is, the upper capacity metric mean dimension of saturated sets carries the full metric mean dimension, as stated in Theorem 1.4.
@@
Let $(X,d,f)$ be a t.d.s. with specification property. If $K\subseteq M_{f}(X)$ is a nonempty compact convex set, and $U\subset X$ is a nonempty open set, then
$\overline{\rm mdim}_{M}^{P}\left(f,G_{K},d\right)=$ $\overline{\rm mdim}_{M}^{P}(f,G_{K}^{\omega,U},d)=\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\sup_{\mu\in K}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi),$
where $\overline{\rm mdim}_{M}^{P}(f,Z,d)$ is the packing upper metric mean dimension of $Z$, $h_{\mu}(f,\xi)$ is the measure-theoretic entropy of $\mu$ w.r.t. $\xi$.
@@
Let $(X,d,f)$ be a t.d.s. with specification property. If $K\subseteq M_{f}(X)$ is a nonempty compact connected set, and $U\subset X$ is a nonempty open set, then
$\overline{\rm mdim}_{M}^{B}\left(f,G_{K},d\right)=$ $\overline{\rm mdim}_{M}^{B}(f,G_{K}^{\omega,U},d)=\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{\mu\in K}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi),$
where $\overline{\rm mdim}_{M}^{B}(f,Z,d)$ is the Bowen upper metric mean dimension of $Z$.
Perhaps, more satisfactory results would be to exchange the order of $\limsup_{\epsilon\to 0}$ and $\sup_{\mu\in K}$ (or $\inf_{\mu\in K}$). However, up to now, the authors remain unaware of whether there exist any counterexamples among infinite entropy systems with the specification property for which the following equalities hold strictly
and
Inspired by the work of [GS21, CPV24], we define two variants of mean Rényi information dimension of invariant measures. Let $K$ be a closed set of $M_{f}(X)$. For any $\mu\in K$, we use $M_{K}(\mu)$ to denote the space of sequences of invariant measures within $K$ which converge to $\mu$ in the weak-topology. We define the (modified) lower and upper mean Rényi information dimensions of $\mu$ as
$\underline{\rm mrid}_{\mu}(f,K,d):$ $=\inf_{(\mu_{\epsilon})_{\epsilon}\in M_{K}(\mu)}\{\liminf_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{\mathrm{diam}(\xi)<\epsilon}h_{\mu_{\epsilon}}(f,\xi)\},$
$\overline{\rm mrid}_{\mu}(f,K,d):$ $=\sup_{(\mu_{\epsilon})_{\epsilon}\in M_{K}(\mu)}\{\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{\mathrm{diam}(\xi)<\epsilon}h_{\mu_{\epsilon}}(f,\xi)\}$
respectively. Then this allows us to obtain the the variational principles in terms of the (modified) lower and upper mean Rényi information dimensions.
@@
Let $(X,d,f)$ be a t.d.s. with specification property. If $K\subseteq M_{f}(X)$ is a nonempty compact convex set, then
$\underline{\rm mdim}_{M}^{B}(f,G_{K},d)=\inf_{\mu\in K}\{\underline{\rm mrid}_{\mu}(f,K,d)\},$
$\overline{\rm mdim}_{M}^{P}(f,G_{K},d)=\sup_{\mu\in K}\{\overline{\rm mrid}_{\mu}(f,K,d)\}.$
We remark that Theorem 1.3 does not require the connectedness of $K$ since any convex set in a topological vector space is automatically connected.
@@
Let $(X,d,f)$ be a t.d.s. with specification property. Let $K\subseteq M_{f}(X)$ be a nonempty compact connected set, and $U\subset X$ be a nonempty open set. Then
@@
As an immediate consequence of Theorems 1.1, 1.2 and 1.4, the most interesting case occurs when $K$ consists of exactly a single point. The Corollary 1.5 extends the result given in [YCZ25, Theorem 1.1] to all non-ergodic measures for systems with specification property.
@@
Let $(X,d,f)$ be a dynamical system with specification property. Then for every $\mu\in M_{f}(X)$,
@@
Consequently, if $\mu\in M_{f}(X)$ is a maximal metric mean dimension measure in the sense of $\overline{\rm mdim}_{M}(f,X,d)=\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi)$, then we have
$\overline{\rm mdim}^{B}_{M}(f,G_{\mu},d)=\overline{\rm mdim}^{P}_{M}(f,G_{\mu},d)=\overline{\rm mdim}_{M}^{UC}(f,G_{\mu},d)=\overline{\rm mdim}_{M}(f,X,d).$
The “Consequently” part can be realized for some dynamical systems. For example, the full shift over the infinite product of $m$-dimensional cubes $[0,1]^{m}$ with the compatible product metric $d(x,y)=\sum_{n\in\mathbb{Z}}\frac{||x_{n}-y_{n}||_{\mathbb{R}^{m}}}{2^{|n|}}$, and $\mu=\mathcal{L}^{\otimes\mathbb{Z}}$ is product measure of the Lebesgue measure $\mathcal{L}$ on $[0,1]^{m}$.
The concept of maximal metric mean dimension measure was proposed in [YCZ23a]. However, determining what types of infinite entropy systems can possess a maximal metric mean dimension measure remains a challenging task. See [LR24] for the existence of maximal metric mean dimension measure for typical conservative homeomorphisms. By Gutman- Śpiewak variational principle (as presented in Theorem 2.4) and Corollary 1.5, we have an abstract criterion for maximal metric mean dimension measure for identifying maximal metric mean dimension measures in systems with the specification property.
Corollary 1.6.
Let $(X,d,f)$ be a t.d.s. with specification property. An invariant measure $\mu\in M_{f}(X)$ is a maximal metric mean dimension measure if and only if $\overline{\rm mdim}^{B}_{M}(f,G_{\mu},d)=\overline{\rm mdim}_{M}(f,X,d).$
Some comments are in order.
  • (1)
    There are plenty of infinite entropy systems with specification property, for instance, the usual full shift over the infinite product of $m$-dimensional cubes $[0,1]^{m}$, and the examples presented in [BR23, Section IV], and our main results can be applied to these systems.
  • (2)
    In Theorem 1.3, we do not know whether the equalities hold for Bowen upper metric mean dimension and packing lower metric mean dimension. The other theorems mentioned above are valid for the corresponding lower metric mean dimensions by considering changing the $\limsup_{\epsilon\to 0}$ into $\liminf_{\epsilon\to 0}$.
  • (3)
    We believe that it is trivial to extend the aforementioned results to metric mean dimensions with potential by making some modifications to the current proofs. The novelty of our work not only gives a proper and full extension of the previous work in [PS07, ZCC12, ZC13, HTW19] to the setting of infinite entropy systems of $\mathbb{Z}$-actions with specification property, but also may be helpful for studying the multifractal analysis of other types of fractal-like sets in infinite entropy systems. It remains unclear whether there are some possible substitutions of the uniform separation property so that all the above results are true for $g$-almost product property, or even for cases beyond $\mathbb{Z}$.
The paper is organized as follows. In section 2, we recall the precise definitions of metric mean dimensions and the associated properties, and two types of measure-theoretic $\epsilon$-entropies, and the specification property. In section 3, we prove the main results. In section 4, we consider some applications of the main results, and apply it to investigate the level sets of Birkhoff average of a continuous potential, Li-Yorke chaos and the metric mean dimensions of generic points of the full shift over compact metric spaces.

### v4

@@
Abstract. From a geometric perspective, we employ metric mean dimension to investigate the set of generic points of invariant measures and saturated sets in infinite entropy systems. For systems with the specification property, we establish certain variational principles for the Bowen and packing metric mean dimensions of saturated sets in terms of Kolmogorov-Sinai $\epsilon$-entropy, and prove that the upper capacity metric mean dimension of saturated sets has full metric mean dimension. Consequently, the Bowen and packing metric mean dimensions of the set of generic points of invariant measures coincide with the mean Rényi information dimension, and the upper capacity metric mean dimension of the set of generic points of invariant measures also has full metric mean dimension.
As applications, for systems with the specification property, we present the qualitative characterization of the metric mean dimensions of level sets, the set of mean Li-Yorke pairs in infinite-entropy systems, and the set of generic points of invariant measures in full shifts over compact metric spaces.
Keywords: multifractal analysis; metric mean dimension; saturated set; generic point; variational principle
@@ 1 Introduction
Let $(X,d,f)$ be a topological dynamical system (TDS for short), where $(X,d)$ is a compact metric space, and $f:X\to X$ is a homeomorphism. Let $M(X)$, $M_{f}(X)$, $M_{f}^{e}(X)$ denote the sets of Borel probability measures on $X$ endowed with the weak-topology, $f$-invariant, and $f$-ergodic Borel probability measures on $X$, respectively.
Quantitative analysis of dynamical systems aims to use entropy-like quantities to understand the dynamics of the abstract dynamical systems. Among these entropy-like quantities, measure-theoretic entropy and topological entropy are two topological invariants to characterize the topological complexity of dynamical systems, which are related by the classical variational principle:
@@
where $h_{top}(f)$ denotes the topological entropy of $X$, and $h_{\mu}(f)$ is the measure-theoretic entropy of $\mu$. To extend the concept of topological entropy to any non-empty subsets, resembling the definition of Hausdorff dimension, Bowen [Bow73] in his pioneering work introduced Bowen topological entropy on subsets. It turns out that the Bowen topological entropy provides a powerful tool for studying the “size” of fractal-like sets and has directly contributed to the development of multifractal analysis in dynamical systems. The multifractal analysis, with its origin dating back to Besicovitch, takes into account the decomposition of the whole phase space into smaller subsets consisting of points with the similar dynamical behaviors. And then, from a geometrical and topological perspectives, we use the fractal dimension and entropy-like quantity to describe the “size” of these subsets. Such a procedure allows us to obtain more information about the dynamics and even recovers the partial dynamics. This phenomenon is also known as multifractal rigidity.
Typically, infinite entropy systems exhibit rather complicated topological complexity. As a result, the concept of topological entropy fails to offer more information about the dynamics of the system. In 2000, Lindenstrauss and Weiss [LW00] introduced the metric mean dimension, which is a dynamical analogue of box dimension. The measure-theoretic counterpart of metric mean dimension, called upper and lower mean Rényi information dimensions, is due to Gutman and Śpiewak[GS21]. The analogous variational principles for metric mean dimension can be found in the references [LT18, GS21, YCZ25, S22]. Given $x\in X$, denote by
$\mathcal{E}_{n}(x):=\frac{1}{n}\sum_{i=0}^{n-1}\delta_{f^{i}x}$
the $n$-th empirical measure of $x$, where $\delta_{x}$ is the Dirac measure at $x$. It is well-known that, in the weak-topology, the limit-point set $V_{f}(x)$ of the sequence $\{\mathcal{E}_{n}(x)\}_{n\geq 1}$ is a non-empty compact connected subset of $M_{f}(X)$ . In particular, if $V_{f}(x)=\{\mu\}$ for some $f$-invariant measure $\mu$, then $x$ is called a generic point of $\mu$. Denote by $G_{\mu}$ the set of all generic points of $\mu$. The Birkhoff ergodic theorem shows that $\mu(G_{\mu})=1$ if and only if $\mu$ is ergodic.
We aim to study the multifractal analysis of the set of generic points in infinite entropy systems. Let us mention some results about the topological entropy of the set of generic points. For every ergodic measure $\mu$, Bowen [Bow73] proved that the Bowen topological entropy of $G_{\mu}$ equals $h_{\mu}(f)$, i.e.,
@@
For the non-ergodic measure $\mu$, one has $\mu(G_{\mu})=0$, and there exist the examples showing that $G_{\mu}=\emptyset$ but $h_{\mu}(f)>0$. Therefore, Bowen’s result fails for non-ergodic measures. In 2007, Pfister and Sullivan [PS07] proved that Bowen’s equality still holds for non-ergodic measures within a more general framework. More generally, given a non-empty closed connected subset $K\subset M_{f}(X)$, the saturated set $G_{K}$ associated with $K$ is defined by
$G_{K}=\{x\in X:V_{f}(x)=K\}.$
By introducing the $g$-almost product property (e.g. $\beta$-shifts), which is strictly weaker than the specification property, and the uniform separation property (e.g. expansive and more generally asymptotically $h$-expansive systems), they established a variational principle for $G_{K}$:
$h_{top}^{B}(T,G_{K})=\inf_{\mu\in K}h_{\mu}(f).$
Furthermore, under only the assumption of the $g$-almost product property, they [PS07] showed that $h_{top}^{B}(f,G_{\mu})=h_{\mu}(f)$ for all $\mu\in M_{f}(X)$. For systems with the specification property and positive expansive property, Zhou, Chen and Cheng [ZCC12] also obtained a (dual) variational principle for packing topological entropy of $G_{K}$:
$h_{top}^{P}(f,G_{K})=\sup_{\mu\in K}h_{\mu}(f).$
Later, for systems with the $g$-almost product property, Hou, Tian and Zhang [HTZ23] proved that the upper capacity topological entropy of the saturated sets with more complicated dynamics has full topological entropy, i.e.,
@@
where $G_{K}^{C}:=G_{K}\cap\{x\in X:C_{f}(X)\subset\omega_{f}(x)\},$ $\omega_{f}(x):=\bigcap_{n=0}^{\infty}\overline{\bigcup_{k=n}^{\infty}\{f^{k}x\}}$ is the $\omega$-limit set of $x$, and $C_{f}(X)$ is the measure center given by $C_{f}(X):=\overline{\cup_{\mu\in M_{f}(X)}{\rm supp}{(\mu)}}$, ${\rm supp}(\mu)$ is the supported set of $\mu$. Then, in the context of the infinite entropy systems, the aforementioned work suggests the following questions for saturated sets:
@@
For (1), the authors in [YCZ25] proved that the packing metric mean dimension of the set of generic points equals the mean Rényi information dimension. However, this equality holds for Bowen metric mean dimension under the assumption of an analogous Brin-Katok formula within the context of infinite entropy systems ([YCZ25, §5, Question,(1)]); for (2), since the uniform separation property implies the system has finite topological entropy, extending the work of [PS07, ZCC12, HTZ23] to the framework of infinite entropy systems encounters some nontrivial challenges. Actually, since the asymptotic behavior of orbits ${\rm orb}_{f}(x)=\{f^{n}(x):n\geq 0\}$ is a fundamental topic in ergodic theory, the subsequent statements of main results involve more general (mixed) saturated sets. For simplicity, we write
@@
where $U\subset X$ is a non-empty open subset, $K\subset M_{f}(X)$ is a non-empty closed connected (or convex) subset. If $G_{K}^{\omega,U}\not$$\emptyset$, then there exists a transitive point in the sufficiently “small” open set $U$ such that $V_{f}(x)=K$. For systems with the specification property, we provide a complete answer to question (2).
Our main results clarify that the saturated sets exhibit different dynamical behaviors in terms of different types of metric mean dimensions. There are certain variational principles for the Bowen and packing metric mean dimensions of saturated sets, as presented in Theorems 1.1, 1.2 and 1.3.
Let $\overline{\rm mdim}_{M}^{B}(f,Z,d)$, $\overline{\rm mdim}_{M}^{P}(f,Z,d)$ and $\overline{\rm mdim}_{M}^{UC}(f,Z,d)$ denote the Bowen/packing/upper capacity metric mean dimension of $Z$, respectively.
@@
Let $(X,d,f)$ be a TDS with the specification property. If $K\subseteq M_{f}(X)$ is a non-empty compact convex set, and $U\subset X$ is a non-empty open set, then
$\overline{\rm mdim}_{M}^{P}\left(f,G_{K},d\right)=$ $\overline{\rm mdim}_{M}^{P}(f,G_{K}^{\omega,U},d)=\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\sup_{\mu\in K}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi).$
@@
Let $(X,d,f)$ be a TDS with the specification property. If $K\subseteq M_{f}(X)$ is a non-empty compact connected set, and $U\subset X$ is a non-empty open set, then
$\overline{\rm mdim}_{M}^{B}\left(f,G_{K},d\right)=$ $\overline{\rm mdim}_{M}^{B}(f,G_{K}^{\omega,U},d)=\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{\mu\in K}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi).$
@@
Let $(X,d,f)$ be a TDS with the specification property. If $K\subseteq M_{f}(X)$ is a non-empty compact convex set, then
$\underline{\rm mdim}_{M}^{B}(f,G_{K},d)=\inf_{\mu\in K}\underline{\rm mrid}_{\mu}^{B}(f,K,d),$
$\overline{\rm mdim}_{M}^{P}(f,G_{K},d)=\sup_{\mu\in K}\overline{\rm mrid}_{\mu}^{P}(f,K,d),$
where $\underline{\rm mrid}_{\mu}^{B}(f,K,d)$ and $\overline{\rm mrid}_{\mu}^{P}(f,K,d)$ denotes the Bowen and packing Rényi information dimensions of $\mu$, respectively.
We remark that any convex set in a topological vector space is automatically connected. Although the $\mu$-measure of $G_{\mu}$ is zero for every non-ergodic measure $\mu$, its topological complexity can be sufficiently large, that is, the upper capacity metric mean dimension of saturated sets carries the full metric mean dimension, as stated in Theorem 1.4.
@@
Let $(X,d,f)$ be a TDS with the specification property. Let $K\subseteq M_{f}(X)$ be a non-empty compact connected set, and $U\subset X$ be a non-empty open set. Then
@@
As an immediate consequence of Theorems 1.1, 1.2 and 1.4, the most interesting case occurs when $K\subset M_{f}(X)$ consists of exactly a single point. The following Corollary 1.5 extends the result [YCZ25, Theorem 1.1] to all non-ergodic measures for systems with the specification property.
@@
Let $(X,d,f)$ be a TDS with the specification property. Then for every $\mu\in M_{f}(X)$,
@@
Except Theorem 1.3, all theorems mentioned above are valid for the corresponding lower metric mean dimensions by considering changing the $\limsup_{\epsilon\to 0}$ into $\liminf_{\epsilon\to 0}$. Besides, it becomes unclear whether the aforementioned results for weaker specification-like property.
The paper is organized as follows. In section 2, we recall the precise definitions of metric mean dimensions and the associated properties, and two types of measure-theoretic $\epsilon$-entropies, and the specification property of TDSs. In section 3, we prove the main results. In section 4, we consider some applications of the main results, and apply it to investigate the level sets of the Birkhoff average of a continuous potential, Li-Yorke chaos in finite entropy systems and the metric mean dimensions of generic points of the full shift over compact metric spaces.